# Self-Regularization with Sparse Autoencoders for Controllable LLM-based Classification

### v2

@@ 3.2.1. Extracting Features from Latent Space with SAEs
where $\mathbf{a}=\sigma(\mathbf{x}\cdot\mathbf{W})$ is the activation vector of the sparse autoencoder $h$ on input $\mathbf{x}$, and $\lambda$ is a hyper-parameter. In practice, we leverage Top-K SAE (Makhzani and Frey, 2013) in our work, which only allows Top-$K$ most activated feature to reconstruct any input representation, explicitly regularizing the sparsity of $h$. $\mathcal{L}_{\text{SAE}}$ indicates that a sparse autoencoder reconstructs arbitrary hidden representation $\mathbf{x}$ by using as few as its learned weight vectors, i.e., $\mathbf{W}^{\top}[c]$ for $c=1,...,C$. The sparse constraint on the activations of $h$ leads to a nice property to understand the latent space, as each dimension $c$ is expected to be activated by a particular pattern, indicating the nature of monosemantic instead of polysemantic. Therefore, each learned weight vector can be considered as an extracted feature. Then, to understand the semantic meaning of each feature, we collect $M$ text spans $\mathcal{I}^{(c)}\subset\mathcal{X}$ whose hidden representations maximally activate a learned feature vector $\mathbf{W}^{\top}[c]$ as the final text-based explanation for the $c$-th feature, i.e.,
@@ 3.2.2. Adapting Pre-trained SAE to Downstream Task
The size of the training dataset $\mathcal{D}$ on one classification task is too small to train sparse autoencoder $h$, which easily leads to the “dead feature” problem (Bricken et al., 2023). One way to overcome this problem is to train the sparse autoencoder on a large and general corpus. However, this method may fail to let the sparse autoencoder capture task-specific features that preserve the information in $\mathcal{D}$. Thus, we propose a two-stage training framework to overcome this challenge as illustrated in Figure 1. In the first stage, we pre-train the sparse autoencoder on a large and general dataset based on Equation 2. In the second stage, we identify dead features that have not been activated on $\mathcal{D}$, denoted as $\widetilde{\mathbf{W}}$. We then fine-tune sparse autoencoder $h$ while encouraging dead features $\widetilde{\mathbf{W}}$ to reconstruct the residual $\mathbf{r}$ of the normal activated features, i.e.,
@@ 4.1. General Settings
We focus on Mistral-7B-Instruct (Jiang et al., 2023) as it has shown strong performance on various downstream tasks compared to other models. Following previous work (Wang et al., 2023b), the last hidden state of input texts on the skip-connect stream is considered as the representation of the texts. We collect the hidden states from the $8$th $16$th, $24$th, and $32$th layers from Mistral-7B, which refers to around a total of $25\%$, $50\%$, $75\%$ and $100\%$ layers are passed as suggested by (Lieberum et al., 2024). Our main experiments are conducted based on the hidden embeddings from the $16$th layer as the vanilla classifier achieves the best performance on this layer (see discussions in Section 4.5.2). The sampling configuration for text generation with 0.7 temperature and 1.0 top-p is set as suggested by (AI, 2024).
@@
We pre-train our SAE following established methods (Gao et al., 2024; Lieberum et al., 2024), using a curated dataset of 711K unique queries spanning diverse instruction-tuning sources (RyokoAI, 2023; Ding et al., 2023; Bai et al., 2022; Liu et al., 2023; Xu et al., 2023; Wang et al., 2024). The Top-K SAE with $K=20$, initialized with $2^{16}$ feature vectors via Kaiming initialization (He et al., 2015), adheres to scaling laws (Gao et al., 2024) and is trained using AdamW (Loshchilov et al., 2017) over five epochs with a constant learning rate $1e^{-3}$ and batch size $512$ instances. We further fine-tune the sparse autoencoders on downstream tasks with a default configuration of $5e^{-5}$ learning rate and $512$ batch size over 5 epochs. Appendix B includes detailed settings about pre-training and fine-tuning sparse autoencoders.
@@ 4.2. RQ1: Effectiveness
To evaluate the effectiveness of the proposed self-regularization framework, we first consider several baselines that utilize the hidden embeddings of LLMs as inputs to improve the generalizability of classifiers, where the early stop strategy is applied as default. w/o Regularization takes the raw hidden embeddings of LLMs as input and is trained over 50 epochs without any regularization strategy. Early Stopping monitors the validation accuracy at the end of each epoch, and the checkpoint with the best validation accuracy is used for evaluation. Dropout (Srivastava et al., 2014) randomly assign $p$of input features to 00 to improve the generalizability of classifier, where we grid search $p\in\{5,10,20,30\}$ and report the best performance. Weight Regularization (Hoerl and Kennard, 1970) applies the $l2$-norm of the classifier weight as a constraint during classifier training, where the hyper-parameter is grid searched from $\{1e^{-2},1e^{-3},1e^{-4}\}$. Focal Loss (Ross and Dollár, 2017) modifies the standard cross-entropy loss by down-weighting the loss contribution of well-classified examples, focusing the classifier on harder, misclassified samples. We tune the hyper-parameter $\gamma\in\{1,2,3\}$. Label Smoothing (Müller et al., 2019) replaces hard one-hot labels with a weighted combination of the ground truth and uniform distribution over all classes, encouraging the model to produce more calibrated predictions. We search the smoothing factor $\alpha\in\{0.05,0.1,0.2\}$ and report the best results. Spectral Regularization (Yoshida and Miyato, 2017) penalizes the high spectral norm of weight matrices in the classifiers, reducing the sensitivity to input perturbation. We manually tune the hyper-parameter in a range of $[1e^{-2},1e^{-4}]$. Note that, the early stop strategy is applied to all baselines, except of “w/o Regularization”. Apart from the above training-based baselines, we also consider several prompting-based methods. Direct Prompt (AI, 2024) prompts LLM (i.e., Mistral-7B-Instruct-v0.2) to make prediction with a human written instruction. CoT Prompt (Wei et al., 2022) further requests the LLM to provide an analysis before making final decisions. The performances of the baselines and our proposed method over multiple random seeds are listed in Table 1.
@@ 4.3. RQ2: Ablation Studies
Pre-trained $29{,}498$ 0.4466 $27{,}796$ 0.5049 $33{,}985$ 0.5012
Fine-tuned $12{,}185$ 0.2288 $15{,}860$ 0.3449 $25{,}056$ 0.3337
@@ 4.5. RQ4: Sensitivity Analysis
Layer=8Layer=16Layer=24Layer=3250505050606060607070707080808080F1 Score (%)Early StoppingSelf-Regularization
@@ Appendix A Minimum Sample Size to Effectively Estimate Mean of Sparse Feature Activations
This result highlights the influence of feature sparsity: as $p$ decreases, the total number of required samples grows proportionally as $1/p$. In practice, if a task-specific feature occurs with a probability $p=0.01$ in the documents (which is already an ideal situation), estimating the mean activation within a margin $d=0.1\times\sigma$ at a 95% confidence level would require approximately:
@@ Appendix B Details of Training Sparse Autoencoders
The pre-training process and hyperparameter configuration adhere to established practices outlined in prior research (Bricken et al., 2023; Gao et al., 2024; Lieberum et al., 2024). For pre-training, we curated data from diverse instruction-tuning datasets, including ShareGPT (RyokoAI, 2023), UltraChat (Ding et al., 2023) (randomly sampling 400,000 instances), HH-RLHF (Bai et al., 2022), WebGLM-QA (Liu et al., 2023), Evol-Instruct (Xu et al., 2023), and HelpSteer2 (Wang et al., 2024), while removing duplicate prompts to produce a total of approximately 711,000 unique queries spanning various topics and intents. This dataset was split into training (90%) and validation (10%) sets, yielding around 113 million tokens for training and 12 million tokens for validation, with an average query length of approximately 178 tokens. We initialize the sparse autoencoder with $2^{16}$ feature vectors using Kaiming initialization (He et al., 2015), selecting this number based on the scaling law, i.e., $C=\mathcal{O}(Z^{\gamma})$, between the feature count $C$ and the training tokens $Z$ observed in (Gao et al., 2024), with $\gamma\approx 0.60$ for GPT2-small, $\gamma\approx 0.65$ for GPT-4, and $\gamma\approx 0.5978$ in our analysis. We also set Top-$K=20$ feature vectors used for each inference following previous work (Gao et al., 2024). We train the sparse autoencoder over five epochs on the pre-training dataset with a batch size of 512 prompts.
@@
We fine-tune the pre-trained sparse autoencoder on the training dataset of each task. At the beginning of each fine-tuning epoch, we use all feature vectors to go through the entire dataset and monitor those feature vectors that have not been activated as dead neurons. The Top-$K=20$ dead feature vectors are asked to reconstruct the residuals of text representations with $\alpha=0.1$. Fine-tuning is performed using the AdamW (Loshchilov et al., 2017) optimizer with $\epsilon=6.25\times 10^{-10}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. For the Toxic Chat Detection and Reward Modeling tasks, fine-tuning is conducted for a total of 5 epochs with a batch size of 512 and a learning rate of $5\times 10^{-5}$. In contrast, for the Disease Diagnosis task, fine-tuning runs for 40 epochs with a batch size of 8 and a learning rate of $3\times 10^{-6}$, as the training dataset is significantly smaller than those of the other two tasks.

### v3

@@ 3.2.1. Extracting Features from Latent Space with SAEs
[…]
@@ 3.2.2. Adapting Pre-trained SAE to Downstream Task
The size of the training dataset $\mathcal{D}$ on one classification task is too small to train sparse autoencoder $h$, which easily leads to the “dead feature” problem (Bricken et al., 2023). One way to overcome this problem is to train the sparse autoencoder on a large and general corpus. However, this method may fail to let the sparse autoencoder capture task-specific features that preserve the information in $\mathcal{D}$. Thus, we propose a two-stage training framework to overcome this challenge as illustrated in Figure 1. In the first stage, we pre-train the sparse autoencoder on a large and general dataset based on Equation 2. In the second stage, we identify dead features that have not been activated on $\mathcal{D}$, denoted as $\widetilde{\mathbf{W}}$. We then fine-tune sparse autoencoder $h$ while encouraging dead features $\widetilde{\mathbf{W}}$ to reconstruct the residual $\mathbf{r}$ of the normal activated features (Gao et al., 2024), i.e.,
@@ 4.1. General Settings
[…]
@@
We pre-train our SAE following established methods (Gao et al., 2024; Lieberum et al., 2024), using a curated dataset of 711K unique queries spanning diverse instruction-tuning sources (RyokoAI, 2023; Ding et al., 2023; Bai et al., 2022; Liu et al., 2023; Xu et al., 2023; Wang et al., 2024). The Top-K SAE with $K=20$, initialized with $2^{16}$ feature vectors via Kaiming initialization (He et al., 2015), adheres to scaling laws (Gao et al., 2024) and is trained using AdamW (Loshchilov et al., 2017) over five epochs with a constant learning rate $1e^{-3}$ and batch size 512512512 instances. We further fine-tune the sparse autoencoders on downstream tasks with a default configuration of $5e^{-5}$ learning rate and 512512512 batch size over 5 epochs. Appendix B includes detailed settings about pre-training and fine-tuning sparse autoencoders.
@@ 4.2. RQ1: Effectiveness
To evaluate the effectiveness of the proposed self-regularization framework, we first consider several baselines that utilize the hidden embeddings of LLMs as inputs to improve the generalizability of classifiers, where the early stop strategy is applied as default. w/o Regularization takes the raw hidden embeddings of LLMs as input and is trained over 50 epochs without any regularization strategy. Early Stopping monitors the validation accuracy at the end of each epoch, and the checkpoint with the best validation accuracy is used for evaluation. Dropout (Srivastava et al., 2014) randomly assign $p$of input features to 0 to improve the generalizability of classifier, where we grid search $p\in\{5,10,20,30\}$ and report the best performance. Weight Regularization (Hoerl and Kennard, 1970) applies the $l2$-norm of the classifier weight as a constraint during classifier training, where the hyper-parameter is grid searched from $\{1e^{-2},1e^{-3},1e^{-4}\}$. Focal Loss (Ross and Dollár, 2017) modifies the standard cross-entropy loss by down-weighting the loss contribution of well-classified examples, focusing the classifier on harder, misclassified samples. We tune the hyper-parameter $\gamma\in\{1,2,3\}$. Label Smoothing (Müller et al., 2019) replaces hard one-hot labels with a weighted combination of the ground truth and uniform distribution over all classes, encouraging the model to produce more calibrated predictions. We search the smoothing factor $\alpha\in\{0.05,0.1,0.2\}$ and report the best results. Spectral Regularization (Yoshida and Miyato, 2017) penalizes the high spectral norm of weight matrices in the classifiers, reducing the sensitivity to input perturbation. We manually tune the hyper-parameter in a range of $[1e^{-2},1e^{-4}]$. Note that, the early stop strategy is applied to all baselines, except of “w/o Regularization”. Apart from the above training-based baselines, we also consider several prompting-based methods. Direct Prompt (AI, 2024) prompts LLM (i.e., Mistral-7B-Instruct-v0.2) to make prediction with a human written instruction. CoT Prompt (Wei et al., 2022) further requests the LLM to provide an analysis before making final decisions. The performances of the baselines and our proposed method over multiple random seeds are listed in Table 1.
@@ 4.3. RQ2: Ablation Studies
Pre-trained 29,49829{,}49829 , 498 0.4466 27,79627{,}79627 , 796 0.5049 33,98533{,}98533 , 985 0.5012
Fine-tuned 12,18512{,}18512 , 185 0.2288 15,86015{,}86015 , 860 0.3449 25,05625{,}05625 , 056 0.3337
@@ 4.5. RQ4: Sensitivity Analysis
Layer=8Layer=16Layer=24Layer=32505050606060707070808080F1 Score (%)Early StoppingSelf-Regularization
@@ Appendix A Minimum Sample Size to Effectively Estimate Mean of Sparse Feature Activations
[…]
@@ Appendix B Details of Training Sparse Autoencoders
[…]
@@
[…]